OpenStudy (liizzyliizz):

not sure if im doing this right for Lim ln(x-1)^2/ ln(x-1) as x->2

OpenStudy (anonymous):

use logarithm properties to get the ^2 out so it is \[(2\times \ln (x-1))\div \ln (x-1)\]

OpenStudy (anonymous):

simplifies to 2

OpenStudy (liizzyliizz):

Are you serious. here i was doing all sorts of crazy things.. when in reality this was a one step problem... UGHHHH